A boy is 26 years younger than his father. In 3 years time, his age would be a third of the father's age. Find the boy's present age.
Answer by mananth(16949)

brother's age x

sons age x
fathers age x+26
3 years later
sons age x+3
fathers age x+26+3 =x+29
if son is 1/3 of father's age , father is 3 times sons age
fathers age 3 times older
x+29=3(x+3)
x+29=3x +9
3x-x=29 -9
2x=20
/2
x=10 sons age
x + 26 = 36 years fathers age